If the heat of combustion for the reaction of Octane is -5471kj/mol, how much heat, in kj, is produced when one gallon of octane is burned in excess oxygen if the reaction works perfectly? Please report your answer in six significant figures.


1
Expert's answer
2020-06-29T07:51:45-0400

1 gallon = 3.785 l

n = 3.785 / 22.4 = 0.1690 mol

Q = 0.1690 × 5471 = 924.5 kJ

Answer: 924.5 kJ.